OpenStudy (anonymous):

HELP ASAP WILL FAN AND MEDAL what are the first four terms of the sequence represented by the expression n(n - 1)- 5? -7, -5, -3, 1 -5, -10, -15, -20 0, 2, 6, 12 -5, -3, 1, 7

OpenStudy (anonymous):

given the function rule f(x)=x^2-4x+3 what is the output of f(-3) 7 -5 25 5

OpenStudy (anonymous):

\[n(n - 1)- 5\] put \(n=1\) get \(1\times (1-1)-5=-5\) for the first term

OpenStudy (anonymous):

for the second, put \(n=2\) get \[2\times (2-1)-5\]

OpenStudy (anonymous):

\[ f(x)=x^2-4x+3\] \[ f(-3)=(-3)^2-4\times (-3)+3\]